OpenStudy (anonymous):

Whats 3 9/10 - 1 2/5???

OpenStudy (idily101):

find the common factor

OpenStudy (anonymous):

3-1 = 2 9/10 - 4/10 = 1/2 answer = 2 1/2

OpenStudy (anonymous):

You want a com mom bottom bumber so. 12/5 would be mulitplied by to so you would have the common number 10. 39/10-24/10=15